OpenStudy (anonymous):

Which expressions are equivalent to: -5/x-3?

OpenStudy (lucaz):

do you have options?

OpenStudy (anonymous):

Yes.

OpenStudy (lucaz):

what are the options?

OpenStudy (anonymous):

It's not letting me draw... hold

OpenStudy (lucaz):

yeah, it's broken

OpenStudy (lucaz):

they are all equivalent =]

OpenStudy (lucaz):

can you see why?

OpenStudy (anonymous):

How...?

OpenStudy (lucaz):

-a / b is the same as a / -b right?

OpenStudy (lucaz):

like -8/2 = -4 and 8/-2 = -4

OpenStudy (lucaz):

@Yovovel ?

OpenStudy (anonymous):

yes, but is there a way to solve these problems? I could see the first one being the same, but. . .

OpenStudy (anonymous):

you could plug in a value for x and see what you get the ones with the same outcome would be equivalent

OpenStudy (lucaz):

multiply -5 by -1 and (x-3) by -1

OpenStudy (lucaz):

or do what helpme1.2 says..

OpenStudy (anonymous):

I'll try

OpenStudy (anonymous):

Oh! It works. I'll try the next question.

OpenStudy (lucaz):

good luck

OpenStudy (anonymous):

Is this right?

OpenStudy (anonymous):

Wait, do I do the same if there's y and x?

OpenStudy (lucaz):

you can do it, just substitute the same value for x's and the same value for y's y(you can pick the same value for both x and y) but I think its easir using algebra for more than 1 variable

OpenStudy (anonymous):

Hmm.. -x-7/y

OpenStudy (lucaz):

I think it's easier multiplying both numerator and denominator by some value to find and equivalent expression

OpenStudy (lucaz):

hey, you can do the same process, just pick a value for x and other for y

OpenStudy (anonymous):

I put an x on the ones that aren't similar and check mark for the right ones, which was only one.

OpenStudy (lucaz):

they look all equivalent again

OpenStudy (anonymous):

The first one isn't, I tried the same method

OpenStudy (lucaz):

let's say x=3 and y=1 -(3+7)/1 = -10 and (-3-7)/1 = -10

OpenStudy (lucaz):

they just factored out -1 from (-x-7)

OpenStudy (lucaz):

the last option is not equivalent..

OpenStudy (lucaz):

just to make it clear -(a/b) = -a/b = a/-b

OpenStudy (lucaz):

the negative sign means -1 times the fraction